OpenStudy (anonymous):

An equation of a circle is x2 + y2 + 10x + 6y + 18 = 0. Part 1: Show all your work in determining the center and radius of this circle. (3 points)

OpenStudy (anonymous):

(x+5)^2+(y+3)^2=4^2 so centre(-5,-3) radius=4

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

(x^2+2.5x+25)+(y^2+2.3x+3^2)+18-25-9=0 (x+5)^2+(y+3)^2=16